OpenStudy (anonymous):

Write the polynomial in standard form and identify the zeros of the function and the x-intercepts of its graph. f(x)=(x-1)(x-1)(x+2i)(x-2i) So far I have foiled and got the polynomial in standard form as : f(x)=x^4-2x^3-5x^2-8x+4 How do I find the zeros?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

your initial product form was: f(x) = (x-1) (x-1) (x+2i) (x-2i) correct?

OpenStudy (anonymous):

Yes.

OpenStudy (amistre64):

and, when you are multiply number together; and one of them is equal to zero; what does that do for the answer?

OpenStudy (amistre64):

for example: 8*6*23*n = ? ; when n=0

OpenStudy (anonymous):

It makes it zero.

OpenStudy (amistre64):

thats the key to this problem then: (x-1) (x-1) (x+2i) (x-2i) = 0 when any of the (...) parts equal zero

OpenStudy (anonymous):

So I plug in numbers for x, and the zeros are whichever numbers, when plugged into the problem, make the equation = 0?

OpenStudy (amistre64):

exactly, therefore (x-1) = 0 when x=? (x-1) = 0 when x=? (x+2i) = 0 when x=? (x-2i) = 0 when x=?

OpenStudy (anonymous):

so my zeros are... 1, and + or - 2i?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Oh, thank you.

OpenStudy (amistre64):

youre welcome, and good luck ;)

OpenStudy (anonymous):

:)